3 times the difference of a number and 8 is 7 less than a number
Answer by Alan3354(69443)

3 times the difference of a number and 8 is 7 less than a number
-----------
n is a number.
difference = |n - 8|
3*|n-8| = n-7
---
Case 1
3n-24 = n-7
2n = 17
n = 8.5
----
Case 2
24-3n = n-7
31 = 4n
n = 7.75
========================
Notice the difference can be either n-8 or 8-n.